Find an equation of the line satisfying the conditions given - answer in standard form.
Parallel to 3x - y = -5 and passing through (-1,0)

Answer by John10(297)

Find the given line slope by solving for y:
y = 3x + 5
The slope is 3.
Because two lines are parallel, the slope is the same.
Apply the slope-point formula to write the line:
y - yo = m(x - xo)
y - 0 = 3( x - (-1))
y = 3x + 3
which is the line you need.
John10:)